# LaDiR: Latent Diffusion Enhances LLMs for Text Reasoning

### v3

@@ 4.2 Puzzle Planning – Countdown
We evaluate the planning ability of our method using Countdown, a combinatorial arithmetic game. Given a set of input numbers, the goal is to reach a target in $[10,100]$ by applying basic operations ${+,,×,÷}\{$$\times$$\div\}$. Solving a problem thus demands decomposing the target into intermediate subgoals and chaining them correctly. For example, given input numbers $\{97,38,3,17\}$ and target 14, one valid solution is: $97-38=59,\>59-17=42,\>42\div 3=14.$ Following Gandhi et al. (2024), we construct a dataset of 500k examples, holding out 10% of target numbers for out-of-distribution evaluation as test set. We study two settings of growing complexity: CD-4 and CD-5, which use four and five input numbers, respectively.
@@
We compare our method against both autoregressive and diffusion-based approaches. For the autoregressive setting, we include (1) LLaMA 8B SFT, which shares the same base model as ours and is finetuned on the same dataset. For diffusion-based baselines, we consider (2) LLaDA 8B SFT (Nie et al., 2025) and (3) Dream 7B Base (Ye et al., 2025b), two diffusion-based general-purpose language models (the latter evaluated without finetuning), as well as (4) MGDM (Ye et al., 2024a), a small task-specific multinomial-guided diffusion model trained for Countdown.
@@ Results
On the Countdown tasks, as shown in Table 2, our method outperforms autoregressive baselines and remains competitive with specialized diffusion models. In CD-4, it improves Pass@1 by more than 25 points over LLaMA 8B SFT and over 20 points over LLaDA SFT, demonstrating stronger planning ability beyond token-by-token generation, while also delivering the best Pass@100 and over two points higher diversity than any baseline. On the more challenging CD-5 task, our model surpasses AR baselines by nearly 30 points in Pass@1 and over 30 points in Pass@100, again with the highest diversity. In addition, as shown in Figure 3, our pass@k curve rises steeply with $k$, surpassing MGDM at larger $k$. This high pass@k reflects both diverse trajectory exploration and strong potential for reinforcement learning for post-training (Yue et al., 2025).
@@ 4.4 Analysis
Iterative Refinement
Table 3 shows how the our flow-matching model refines its reasoning across denoising steps. From pure noise at $T=1$, the model quickly produces structured equations, though early steps contain arithmetic errors (e.g., off-by-one mistakes). As denoising progresses, partial results stabilize—such as $43+9=52$ appearing consistently from $T=0$ onward—and later steps are gradually corrected. By $T=0.25$, the full reasoning matches the ground truth and remains stable through $T=0$. This demonstrates that our method exhibits the same iterative refinement ability as reasoning models (Shao et al., 2024), progressively correcting previously generated steps. See Table 6 for an example on GSM8K.
@@
We provide additional analyses on interpretability and semantic-level reasoning in Appendix D.
@@ 5 Related Works
Latent reasoning methods address token-level limits of chain-of-thought by enabling reasoning in a latent space, yielding more abstract representations through discrete special tokens that expand internal computation or capture unstated steps (Herel & Mikolov, 2024; Pfau et al., 2024; Wang et al., 2024; Zelikman et al., 2024; Jin et al., 2025). Prior works show that reasoning in latent space, rather than discrete tokens, improves performance by allowing LLMs to generate continuous tokens, either self-generated or provided by an auxiliary model (Cheng & Durme, 2024; Hao et al., 2024; Liu et al., 2024b; Shen et al., 2025; Tack et al., 2025; Zhu et al., 2025; Butt et al., 2025; Zhang et al., 2025; Wu et al., 2025). This has been further extended to recurrent or looped architectures that induce latent reasoning internally, removing the need to represent reasoning steps explicitly as tokens (Chen et al., 2025c; Geiping et al., 2025; Mohtashami et al., 2025; Saunshi et al., 2025; Yu et al., 2025b). However, prior latent reasoning approaches lack interpretability, as their continuous states are opaque and difficult to understand or control. whereas our method structures the latent space with a VAE, making each step explicit and thus more transparent. Due to the page limit, we discuss further related works in Appendix A.
@@ 6 Conclusion
Acknowlegement
We thank Apple for their support and sponsorship of this project, as well as for the insightful discussions and constructive feedback provided during the meetings. This work is supported in part by the NSF Award CCF-2112665 (TILOS), the DARPA AIE program, and the CDC-RFA-FT-23-0069. We thank Vincent Tu for early contributions to this project.
@@ Appendix A Additional Related Works
Masked diffusion language models attempt to address some common limitations of autoregressive LLMs—such as rigid left-to-right decoding and inefficiency—by iteratively denoising masked tokens, enabling parallel and order-agnostic text generation. Prior studies show that these models achieve better inference efficiency compared to AR models while maintaining comparable performance on both general tasks (Zheng et al., 2024; Gong et al., 2025; Nie et al., 2025; Shi et al., 2025a; Song et al., 2025; Ye et al., 2025c; b) and reasoning benchmarks with chain-of-thought (Gao et al., 2024; Ye et al., 2024b; van Krieken et al., 2025; Ye et al., 2025a). However, these approaches remain constrained to language space, unable to capture reasoning at an abstract semantic level or revise previously generated tokens as continuous diffusion models (Ho et al., 2020; Song et al., 2020) do, and they require training on massive datasets rather than leveraging a well-trained LLM. In contrast, our method overcomes these limitations by structuring reasoning in an interpretable continuous latent space, producing abstract CoT representations with self-correction ability for an existing LLM, while keeping diffusion’s strengths in parallel generation to enhance exploration and diversity.
@@
Chain-of-thought reasoning refers to methods which elicit LLMs to generate intermediate reasoning steps in language prior to outputting a final answer in order to improve performance on reasoning tasks. This can be accomplished via prompting methods  (Nye et al., 2021; Wei et al., 2023; Khot et al., 2023; Zhou et al., 2023) or through training LLMs (by SFT, RL, or a combination of the two) to output the intermediate reasoning steps  (Yu et al., 2023; Shao et al., 2024; Yu et al., 2025a). Works have also extended CoT to allow LLMs to mimic various tree search algorithms such as BFS or MCTS, which especially improves performance on more complex tasks  (Xie et al., 2023; Yao et al., 2023; Zhang et al., 2024; Bi et al., 2025). Beyond following specific algorithms, works that implement long chain-of-thought (combining extensive reasoning, exploration, and reflection) have also demonstrated improved reasoning performance  (Shinn et al., 2023; Gandhi et al., 2025; Saha et al., 2025; Xie et al., 2025). One overarching limiting factor with these CoT methods is that they fundamentally work at a next-token-prediction level, constraining the outputs to the token space and limiting the model’s horizon.
@@
Other AR-Diffusion hybrid models have shown successful results in rivaling their AR and diffusion counterparts, particularly in multimodal generation and image understanding. The Transfusion  (Zhou et al., 2024a) architecture demonstrated that hybrid models could outperform standard AR models and compete with state-of-the-art diffusion models in image-generation benchmarks, a phenomenon further reinforced by other studies of hybrid models  (Fan et al., 2024; Tang et al., 2024a; Xiao et al., 2024). This extends beyond image generation, with several works demonstrating the effectiveness of hybrid AR-diffusion models in other domains such as image understanding, video generation, and robot control  (Black et al., 2024; Tong et al., 2024a; Chen et al., 2025a; b). Furthermore–similar to our model architecture–works have demonstrated successful adaptations of frozen models for these hybrid AR-diffusion archictures in multimodal domains  (Pan et al., 2025; Shi et al., 2025b). Aside from the difference in domain from these works, many do not use block diffusion for variable-length generations as in LaDiR and we critically introduce CE loss to guide better latent predictions.
@@ B.1 Variational Autoencoder and $\beta$-VAE
The Variational Autoencoder (VAE) (Kingma & Welling, 2013) is a latent-variable model that learns a compressed representation of data $x$ through an encoder–decoder pair. The encoder $q_{\phi}(z|x)$ maps inputs into a distribution over latent variables $z$, typically parameterized as a diagonal Gaussian. The decoder $p_{\theta}(x|z)$ reconstructs the input from $z$, enabling generative sampling. Training maximizes the evidence lower bound (ELBO):
@@
The $\beta$-VAE (Higgins et al., 2017) introduces a hyperparameter $\beta$ to control the KL weight:
@@
Latent diffusion models (LDMs) (Rombach et al., 2022) operate not on raw high-dimensional inputs (e.g., images or sequences), but in a compressed latent space learned by a VAE. This design provides three key advantages:
@@ B.3.1 $\epsilon$-Prediction
where $z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon$. This is the standard DDPM formulation (Ho et al., 2020). It is stable but sometimes less efficient for long horizons.
@@ B.3.2 $x_{0}$-Prediction (DDIM-$x_{0}$)
This corresponds to the DDIM formulation (Song et al., 2020), enabling deterministic sampling and fewer inference steps, but can overfit to data scale.
@@ B.3.3 $v$-Prediction
Proposed by Salimans & Ho (2022), $v$ is defined as a linear combination of noise and clean latent:
@@ Appendix C Additional Ablation Studies
To assess the impact of different training objectives for latent prediction, we compare several widely used formulations. The first baseline is MSE loss, which directly minimizes the mean-squared error between predicted and ground-truth latents but yields the weakest results. We then adopt three DDIM-based (Song et al., 2020) objectives: predicting the clean latent state ($x_{0}$), the added noise vector ($\epsilon$), and the velocity ($v$). These diffusion objectives consistently improve accuracy, highlighting the advantage of explicitly modeling the denoising process rather than relying on direct predictions. Moreover, the flow matching ($u$) objective achieves the strongest gains, suggesting that learning the latent vector field is more effective for capturing the pattern of reasoning compared to DDIM’s denoising objectives.
@@
Appendix D Additional Analysis
@@ E.1 Math Reasoning
Table 8: Summary of datasets used in our experiments. We use DART-MATH (Tong et al., 2024b), MATH (Hendrycks et al., 2021), GSM8K (Cobbe et al., 2021), College-Math (Tang et al., 2024b), DeepMind-Math (Saxton et al., 2019), OlympiaBench-Math (He et al., 2024), TheoremQA (Chen et al., 2023), and Fresh-Gaokao-Math-2023 (Tang et al., 2024b).

### v4

@@ 4.2 Puzzle Planning – Countdown
We evaluate the planning ability of our method using Countdown, a combinatorial arithmetic game. Given a set of input numbers, the goal is to reach a target in $[10,100]$ by applying basic operations ${+,,×,÷}\{$$\times$$\div\}$. Solving a problem thus demands decomposing the target into intermediate subgoals and chaining them correctly. For example, given input numbers $\{97,38,3,17\}$ and target 14, one valid solution is: $97-38=59,\>59-17=42,\>42\div 3=14.$ Following gandhi2024stream, we construct a dataset of 500k examples, holding out 10% of target numbers for out-of-distribution evaluation as test set. We study two settings of growing complexity: CD-4 and CD-5, which use four and five input numbers, respectively.
@@
We compare our method against both autoregressive and diffusion-based approaches. For the autoregressive setting, we include (1) LLaMA 8B SFT, which shares the same base model as ours and is finetuned on the same dataset. For diffusion-based baselines, we consider (2) LLaDA 8B SFT (nie2025largelanguagediffusionmodels) and (3) Dream 7B Base (ye2025dream), two diffusion-based general-purpose language models (the latter evaluated without finetuning), as well as (4) MGDM (ye2024beyond), a small task-specific multinomial-guided diffusion model trained for Countdown.
@@ Results
On the Countdown tasks, as shown in Table 2, our method outperforms autoregressive baselines and remains competitive with specialized diffusion models. In CD-4, it improves Pass@1 by more than 25 points over LLaMA 8B SFT and over 20 points over LLaDA SFT, demonstrating stronger planning ability beyond token-by-token generation, while also delivering the best Pass@100 and over two points higher diversity than any baseline. On the more challenging CD-5 task, our model surpasses AR baselines by nearly 30 points in Pass@1 and over 30 points in Pass@100, again with the highest diversity. In addition, as shown in Figure 3, our pass@k curve rises steeply with $k$, surpassing MGDM at larger $k$. This high pass@k reflects both diverse trajectory exploration and strong potential for reinforcement learning for post-training (yue2025does).
@@ 4.4 Analysis
Iterative Refinement
Table 3 shows how the our flow-matching model refines its reasoning across denoising steps. From pure noise at $T=1$, the model quickly produces structured equations, though early steps contain arithmetic errors (e.g., off-by-one mistakes). As denoising progresses, partial results stabilize—such as $43+9=52$ appearing consistently from $T=0$ onward—and later steps are gradually corrected. By $T=0.25$, the full reasoning matches the ground truth and remains stable through $T=0$. This demonstrates that our method exhibits the same iterative refinement ability as reasoning models (shao2024deepseekmathpushinglimitsmathematical), progressively correcting previously generated steps. See Table 10 for an example on GSM8K.
@@
Inference Efficiency
Table 4 compares inference latency on the MATH dataset using identical hardware and batch size (8). LaDiR with 10 diffusion steps matches the latency of the AR baseline while achieving comparable Pass@1 and higher Pass@100. With 30 steps, LaDiR offers a flexible accuracy-compute trade-off. This efficiency stems from our compact latent representation: each latent block contains only 4 latent tokens, representing on average  22 text tokens, reducing per-step computation and context length compared to autoregressive decoding over long text sequences. We provide additional analyses on interpretability and semantic-level reasoning in Appendix D.
Method Latency Pass@1 Pass@100
LLaMA CoT SFT 4.7 43.1 89.0
LaDiR (10 steps) 4.9 42.9 90.7
LaDiR (30 steps) 14.8 44.9 92.8
Table 4: Inference latency (seconds/example) comparison on MATH dataset.
@@ 5 Related Works
Latent reasoning methods address token-level limits of chain-of-thought by enabling reasoning in a latent space, yielding more abstract representations through discrete special tokens that expand internal computation or capture unstated steps (herel2024thinkingtokenslanguagemodeling; pfau2024letsthinkdotdot; wang2024guiding; zelikman2024quietstar; jin2025disentanglingmemoryreasoningability). Prior works show that reasoning in latent space, rather than discrete tokens, improves performance by allowing LLMs to generate continuous tokens, either self-generated or provided by an auxiliary model (cheng2024compressedchainthoughtefficient; hao2024traininglargelanguagemodels; liu2024expeditingelevatinglargelanguage; shen2025codicompressingchainofthoughtcontinuous; tack2025llmpretrainingcontinuousconcepts; zhu2025reasoning; butt2025soft; zhang2025soft; wu2025llms). This has been further extended to recurrent or looped architectures that induce latent reasoning internally, removing the need to represent reasoning steps explicitly as tokens (chen2025innerthinkingtransformerleveraging; geiping2025scalingtesttimecomputelatent; mohtashami2025cotformer; saunshi2025reasoninglatentthoughtspower; yu2025enhancingautoregressivechainofthoughtloopaligned). However, prior latent reasoning approaches lack interpretability, as their continuous states are opaque and difficult to understand or control. whereas our method structures the latent space with a VAE, making each step explicit and thus more transparent.
Latent Diffusion for Language Generation
Generative text modeling has recently expanded from autoregressive paradigms to diffusion-based approaches that allow for global iterative refinement. One of the first, Diffusion-LM (li2022diffusion), frames generation as denoising continuous word embeddings to enable fine-grained control, a concept  lovelace2023latent; lovelacestop extended by performing diffusion in a compressed latent space for improved quality and diverse generation modes. For sequence-to-sequence tasks, DiffuSeq (gong2022diffuseq) enables parallel generation with high diversity, while PLANNER (zhang2023planner) addresses long-form text by combining a latent semantic diffusion planner with an autoregressive decoder to reduce repetition. Similarly, Cosmos (meshchaninov2025compressed) learns a compressed latent space for diffusion, enabling parallel text generation with robust semantic grounding. In specialized domains, DiffusionDialog (xiang2024diffusiondialog) utilizes latent variables to handle open-ended conversations, whereas CodeFusion (singh2023codefusion) and TreeDiff (zeng2025treediff) apply diffusion to code synthesis. While prior latent diffusion models focus on text generation, they lack the granularity to model the multi-step causal dependencies required for reasoning tasks. We address this by introducing blockwise variable-length diffusion and rollout training, explicitly shifting the objective from generating fluent text to optimizing reasoning trajectories that lead to correct answers. Due to the page limit, we discuss further related works in Appendix A.
@@ Appendix A Additional Related Works
Masked diffusion language models attempt to address some common limitations of autoregressive LLMs—such as rigid left-to-right decoding and inefficiency—by iteratively denoising masked tokens, enabling parallel and order-agnostic text generation. Prior studies show that these models achieve better inference efficiency compared to AR models while maintaining comparable performance on both general tasks (zheng2024reparameterizeddiscretediffusionmodel; gong2025scalingdiffusionlanguagemodels; nie2025largelanguagediffusionmodels; shi2025simplifiedgeneralizedmaskeddiffusion; song2025seeddiffusionlargescalediffusion; ye2025diffusionlanguagemodelsperform; ye2025dream) and reasoning benchmarks with chain-of-thought (gao2024diffucometcontextualcommonsenseknowledge; ye2024diffusionthoughtschainofthoughtreasoning; vankrieken2025neurosymbolicdiffusionmodels; ye2025autoregressiondiscretediffusioncomplex). However, these approaches remain constrained to language space, unable to capture reasoning at an abstract semantic level or revise previously generated tokens as continuous diffusion models (ho2020denoising; song2020denoising) do, and they require training on massive datasets rather than leveraging a well-trained LLM. In contrast, our method overcomes these limitations by structuring reasoning in an interpretable continuous latent space, producing abstract CoT representations with self-correction ability for an existing LLM, while keeping diffusion’s strengths in parallel generation to enhance exploration and diversity.
@@
Chain-of-thought reasoning refers to methods which elicit LLMs to generate intermediate reasoning steps in language prior to outputting a final answer in order to improve performance on reasoning tasks. This can be accomplished via prompting methods  (nye2021workscratchpadsintermediatecomputation; wei2023chainofthoughtpromptingelicitsreasoning; khot2023decomposedpromptingmodularapproach; zhou2023leasttomostpromptingenablescomplex) or through training LLMs (by SFT, RL, or a combination of the two) to output the intermediate reasoning steps  (yu2023metamath; shao2024deepseekmathpushinglimitsmathematical; yu2025flowreasoningtrainingllms). Works have also extended CoT to allow LLMs to mimic various tree search algorithms such as BFS or MCTS, which especially improves performance on more complex tasks  (xie2023selfevaluationguidedbeamsearch; yao2023treethoughtsdeliberateproblem; zhang2024restmcts; bi2025forestofthoughtscalingtesttimecompute). Beyond following specific algorithms, works that implement long chain-of-thought (combining extensive reasoning, exploration, and reflection) have also demonstrated improved reasoning performance  (shinn2023reflexionlanguageagentsverbal; gandhi2025cognitivebehaviorsenableselfimproving; saha2025learningplanreason; xie2025logicrlunleashingllmreasoning). One overarching limiting factor with these CoT methods is that they fundamentally work at a next-token-prediction level, constraining the outputs to the token space and limiting the model’s horizon.
@@
Other AR-Diffusion hybrid models have shown successful results in rivaling their AR and diffusion counterparts, particularly in multimodal generation and image understanding. The Transfusion  (zhou2024transfusionpredicttokendiffuse) architecture demonstrated that hybrid models could outperform standard AR models and compete with state-of-the-art diffusion models in image-generation benchmarks, a phenomenon further reinforced by other studies of hybrid models  (fan2024fluidscalingautoregressivetexttoimage; tang2024hartefficientvisualgeneration; xiao2024omnigenunifiedimagegeneration). This extends beyond image generation, with several works demonstrating the effectiveness of hybrid AR-diffusion models in other domains such as image understanding, video generation, and robot control  (black2024pi0visionlanguageactionflowmodel; tong2024metamorphmultimodalunderstandinggeneration; chen2025gokuflowbasedvideo; chen2025janusprounifiedmultimodalunderstanding). Furthermore–similar to our model architecture–works have demonstrated successful adaptations of frozen models for these hybrid AR-diffusion archictures in multimodal domains  (pan2025transfermodalitiesmetaqueries; shi2025lmfusionadaptingpretrainedlanguage). Aside from the difference in domain from these works, many do not use block diffusion for variable-length generations as in LaDiR and we critically introduce CE loss to guide better latent predictions.
@@ B.1 Variational Autoencoder and $\beta$-VAE
The Variational Autoencoder (VAE) (kingma2013auto) is a latent-variable model that learns a compressed representation of data $x$ through an encoder–decoder pair. The encoder $q_{\phi}(z|x)$ maps inputs into a distribution over latent variables $z$, typically parameterized as a diagonal Gaussian. The decoder $p_{\theta}(x|z)$ reconstructs the input from $z$, enabling generative sampling. Training maximizes the evidence lower bound (ELBO):
@@
The $\beta$-VAE (higgins2017beta) introduces a hyperparameter $\beta$ to control the KL weight:
@@
Latent diffusion models (LDMs) (rombach2022high) operate not on raw high-dimensional inputs (e.g., images or sequences), but in a compressed latent space learned by a VAE. This design provides three key advantages:
@@ B.3.1 $\epsilon$-Prediction
where $z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\,\epsilon$. This is the standard DDPM formulation (ho2020denoising). It is stable but sometimes less efficient for long horizons.
@@ B.3.2 $x_{0}$-Prediction (DDIM-$x_{0}$)
This corresponds to the DDIM formulation (song2020denoising), enabling deterministic sampling and fewer inference steps, but can overfit to data scale.
@@ B.3.3 $v$-Prediction
Proposed by salimans2022progressive, $v$ is defined as a linear combination of noise and clean latent:
@@ Appendix C Additional Ablation Studies
To assess the impact of different training objectives for latent prediction, we compare several widely used formulations. The first baseline is MSE loss, which directly minimizes the mean-squared error between predicted and ground-truth latents but yields the weakest results. We then adopt three DDIM-based (song2020denoising) objectives: predicting the clean latent state ($x_{0}$), the added noise vector ($\epsilon$), and the velocity ($v$). These diffusion objectives consistently improve accuracy, highlighting the advantage of explicitly modeling the denoising process rather than relying on direct predictions. Moreover, the flow matching ($u$) objective achieves the strongest gains, suggesting that learning the latent vector field is more effective for capturing the pattern of reasoning compared to DDIM’s denoising objectives.
@@
VAE Robustness Augmentations.
When training our VAE, to improve the robustness of the VAE latent space, we apply two augmentation strategies during VAE training: (1) adding Gaussian noise to latent representations with standard deviation $k$, and (2) randomly substituting input tokens with probability $p$. Table 6 shows the impact of these augmentations on GSM8K accuracy. Best performance is achieved at $k=3$ and $p=0.3$. Too little augmentation ($k=0$ or $p=0$) results in overfitting to clean inputs, while excessive augmentation ($k=5$ or $p>0.5$) degrades the latent space quality.
Latent Gaussian Noise (p=0.3) Token Substitution (k=3)
$k$ (std) GSM8K Acc (%) $p$ (prob.) GSM8K Acc (%)
0 68.3 0.0 70.2
1 73.4 0.1 78.3
3 84.2 0.3 84.2
5 79.4 0.5 64.0
0.7 32.4
Table 6: Ablation study on VAE robustness augmentations on GSM8K.
Blockization Strategy
We investigate the sensitivity of the model to different blockization strategies by varying the number of sentences per block. Table 7 shows results for 1, 2, and 3 sentences per block. Using more sentences per block requires more latent tokens to maintain reconstruction quality and significantly increases difficulty for the diffusion model. We find that 1 sentence per block with 4 latent tokens offers the best balance between latent compactness and reasoning accuracy.
# Sentences # Latent Tokens GSM8K MATH
1 4 84.2 45.2
2 8 78.4 39.6
3 12 72.0 36.1
Table 7: Ablation study on blockization strategy (sentences per block).
Appendix D Additional Results and Analysis
Pass@k Results on Math Reasoning
Method MATH GSM8K Gaokao DM College Olympia TheoremQA Avg.
Diffusion Language Models — LLaDA 8B
Base Model
CoT SFT 59.9 92.4 43.0 50.7 58.3 10.2 26.3 48.7
Autoregressive Models — LLaMA 3.1 8B
Sol-Only SFT
CoT SFT 49.8 89.0 37.9 52.0 54.9 12.9 25.0 45.9
iCoT
Pause Token
Coconut 39.3 74.3 29.3 36.9 42.9 6.3 14.9 34.8
Discrete Latent 47.3 88.6 39.7 49.5 53.7 17.8 28.5 46.4
LaDiR 63.7 93.7 45.8 54.2 60.3 15.3 30.7 52.0
Table 8: Pass@100 accuracy across in-domain and out-of-domain math benchmarks.
@@ E.1 Math Reasoning
Table 12: Summary of datasets used in our experiments. We use DART-MATH (tong2024dart), MATH (hendrycks2021measuring), GSM8K (cobbe2021training), College-Math (tang2024mathscale), DeepMind-Math (saxton2019analysing), OlympiaBench-Math (he2024olympiadbench), TheoremQA (chen2023theoremqa), and Fresh-Gaokao-Math-2023 (tang2024mathscale).
E.2 Hyperparameters
Table 13 provides a complete summary of all hyperparameters used in our experiments, covering VAE pretraining, Stage-1 teacher-forcing training, Stage-2 rollout training, and inference settings.
Component Hyperparameter Value
VAE Pretraining Latent dimension ($d_{z}$) 512
# latent tokens per block 4
KL weight $\beta$ $1\times 10^{-5}$
Learning rate $2\times 10^{-5}$
Batch size 128
# of Epochs 2
Stage-1 Teacher-Forcing Flow-matching loss weight ($\lambda_{\mathrm{FM}}$) 5
CE loss weight ($\lambda_{\mathrm{Ans}}$) 1
Special-token loss weight ($\lambda_{\mathrm{Spec}}$) 1
Learning rate $1\times 10^{-5}$
Batch size 64
# of Epochs 20
Stage-2 Rollout Training Learning rate $1\times 10^{-5}$
Batch size 12
# of Epochs 20
Inference Classifier-free guidance scale 4
Answer token decoding temperature 0.7
Table 13: Complete hyperparameter settings for all training and inference stages.